OpenStudy (anonymous):

when the price of gasoline dropped $0.05 per gallon, it cost mr.lee $22.68 to fill his 14-gallon gas tank. by what percent did the price per gallon decrease

OpenStudy (anonymous):

let the old price of gas per galon be x . so (x-0.05)*14 =22.68 so x =1.67 . then (0.05/1.67)*100=2.994% is the percentage that the price decreases.

OpenStudy (anonymous):

how did you get 1.67 as x

OpenStudy (anonymous):

imagine that the old price is x and the new one (after decreasion) is x-0.05 $/galon that means if you buy 1 galon you should pay (x-0.05) $ .now you'v bought 14 galon so you should pay 14*(x-0.05) which is equal 22.68 $. then x=(22.68/14)+0.05

OpenStudy (anonymous):

so it will be 3%

OpenStudy (anonymous):

becuase i have multiple choice and the closest to that is 3%

OpenStudy (anonymous):

yes, the last result is (0.05/x)*100 which is 5/1.67=2.994

OpenStudy (anonymous):

ok thanks